# Relatively big projective modules and their applications to direct sum decompositions

### v1

@@ Introduction
For a long time, the study of projective modules was focused on the finitely generated ones. Outside the finitely generated setting, the classical theorem of Kaplansky [kaplansky] stating that every projective module is a direct sum of countably generated projective modules shows that one can just consider the countably generated case. Interesting monoids associated to infinitely generated projective modules have recently been studied in several different contexts: the monoid $V(\Lambda)\sqcup B(\Lambda)$ of relatively big projectives defined, formally, in [wiegand] and that is the main object of interest in this paper, but also the analogous of the Cuntz monoid for general rings defined in [AABPV], and the $\kappa$-monoids defined in [nazemian_smertnig].
@@
The concept of relatively big projective module was introduced by Příhoda in [P2], extending Bass’ uniformly big projectives [bass]. A projective right $\Lambda$-module $P$ is said to be $J$-big, for $J$ an ideal of $\Lambda$, if every countably generated projective $\Lambda$-module whose trace is contained in $J$ is a homomorphic image of $P$. Every $J$-big projective module $P$ is determined by the ideal $J$ and the isomorphism class of $P/PJ$ (cf. Lemma 1.14). A countably generated projective module $P$ is said to be relatively big if there exist $J$ such that $P$ is $J$-big and $P/PJ$ is finitely generated.
@@
To capture the properties of relatively big projectives, we introduced in [wiegand] the monoid $V(\Lambda)\sqcup B(\Lambda)$ of isomorphism classes of relatively big projective modules. This monoid embeds in $V^{*}(\Lambda)$ and can be thought of as the disjoint union of $V(\Lambda/I)$, where $I$ runs through the trace ideals of the countably generated $\Lambda$-modules, see Corollary 2.7.
@@
In [P2], it was proved that, over a noetherian ring $\Lambda$ which satisfies a certain descending condition on two-sided ideals, that we call the ()(*)( ∗ )-condition, $V(\Lambda)\sqcup B(\Lambda)=V^{*}(\Lambda)$ (cf. [P2, Lemma 2.8] and Theorem 1.18). Commutative noetherian rings, finitely generated algebras over commutative noetherian rings of Krull dimension one and noetherian semilocal rings satisfy the ()(*)( ∗ ) condition, therefore all countably generated projective modules over these rings are relatively big. These results were the main ingredient to make a detailed study of projective modules over noetherian semilocal rings in [crelle] or to study some infinite direct sum decompositions in [wiegand].
@@
The main results of this paper give new classes of rings for which $V(\Lambda)\sqcup B(\Lambda)=V^{*}(\Lambda)$. We prove that right noetherian PI-rings satisfy the ()(*)( ∗ )-condition (see Theorem 4.8), and this is one of the main ingredients to show all countably generated right (or left) projective modules over a right noetherian PI-ring are relatively big in Corollary 4.9. In particular, this applies to module-finite algebras over commutative noetherian rings, such as group algebras of finite groups over commutative noetherian rings or to endomorphism rings of finitely generated modules over commutative noetherian rings, so we show that the restriction on Krull dimension we were imposing in our previous works was superfluous.
@@
Outside the noetherian setting, it is easy to show that for semiperfect rings all countably generated projective right (or left) modules are relatively big (cf. Proposition 6.3). In Theorem 6.16, we show that the same conclusion holds for $\Lambda$ a locally semiperfect torsion-free algebra, over an $h$-local domain $R$ with field of fractions $Q$, such that $\Lambda\otimes_{R}Q$ is simple artinian.
If $\Lambda$ is the endomorphism ring of a finitely generated torsion-free module $M$ over an $h$-local domain $R$ then $\Lambda\otimes_{R}Q$ is simple artinian. If, in addition, $R$ has Krull dimension $1$ and all the direct summands of a direct sum of copies of $R\oplus M$ are again a direct sum of finitely generated modules, it was proved in [AHP] that $\Lambda$ is locally semiperfect. The converse of this result is not true, in general, as additional conditions in the ranks of indecomposable modules over the localizations of $R$ at its maximal ideals must be satisfied. Our results on projective modules allow us to prove that such rank conditions are not only necessary but also sufficient to prove the converse (cf. Theorem 9.3). This result complements and clarifies the results in [AHP].
It is interesting to recall that Procesi and Small [PS] proved that the endomorphism ring of a finitely generated module over a commutative ring or, in general, over a ring satisfying a polynomial identity, is also a ring satisfying a polynomial identity. As far as we know, this point of view has not been so much examined in the world of finitely generated modules, endomorphism rings and direct sum decompositions and, certainly, it seems it should be worthwhile exploring it.
The paper is structured as follows. In Section 1 we give the basic definitions and notation. We introduce the needed background, presenting the concept of relatively big projective, the ()(*)( ∗ )-condition and the main results relating both notions. Section 2 we examine the monoid of isomorphism classes of relatively big projective modules, its algebraic properties and explore the right/left (non)symmetry of the monoids; check Example 2.10 and Corollary 2.12. In Section 3 we show that countably generated, projective right modules lift modulo a right $T$-nilpotent ideal, cf. Lemma 3.1. We prove that for a right $T$-nilpotent ideal $N$ of $\Lambda$ there are monoid isomorphisms between $V^{*}(\Lambda)$ and $V^{*}(\Lambda/N)$. This isomorphism restricts to isomorphisms between $V(\Lambda)$ and $V(\Lambda/N)$, and between $V(\Lambda)\sqcup B(\Lambda)$ and $V(\Lambda/N)\sqcup B(\Lambda/N)$, cf. Corollary 3.3. In Section 4, these results are an important ingredient to prove that right noetherian PI-rings satisfy the ()(*)( ∗ )-condition (Theorem 4.8), and to conclude that $V(\Lambda)\sqcup B(\Lambda)=V^{*}(\Lambda)$ for right noetherian PI-rings in Corollary 4.9. In particular, this holds for module-finite algebras over commutative noetherian rings (cf. Corollary 4.10).
@@
In the rest of the paper, we turn to the study of algebras over $h$-local domains. These are domains satisfying strong noetherianity conditions on the maximal spectrum, more precisely, any non-zero ideal of $R$ is contained only in a finite number of maximal ideals, and any non-zero prime ideal of $R$ is contained in a unique maximal ideal. In particular, noetherian domains of Krull dimension one are always $h$-local. The definition of $h$-local domain is attributed to Matlis. However, Jaffard in [jaffard] introduced Dedekind-type domains as those domains $R$ such that any nonzero ideal is a finite product of ideals that are contained in a unique maximal ideal of $R$. Jaffard proved that Dedekind-type domains coincide with $h$-local domains, cf. [FHL, Theorem 2.1.5].
@@
The class of $h$-local domains is ubiquitously associated with many classical problems in the study of modules over non-noetherian domains, see [matlis3] and the recent survey [olberding]. Non-finitely generated projective modules over an $h$-local domain are free [hinohara], so what is of interest is the description of projective modules on algebras over such domains.
Our main motivation to study this setting comes from our paper [AHP], in which we studied domains for which the class of direct sums of finitely generated torsion-free modules is closed under direct summands. The type of algebras over $h$-local domains we consider appears naturally when considering this question. We will find some neat applications to this question of the theory we develop in the final section of the paper.
@@
Section 5 discusses Package Deal Theorems for projective modules over $\Lambda$, whenever $\Lambda$ is an algebra over an $h$-local domain; our goal is twofold, recalling the results from [AHP] and proving some extension needed in the subsequent sections. Section 6 focuses on semiperfect and locally semiperfect algebras. We observe in Proposition 6.3 that $V(\Lambda)\sqcup B(\Lambda)=V^{*}(\Lambda)$ for semiperfect rings.
@@
Let $R$ be an $h$ local domain with field of fractions $Q$. In the final part of the paper, we study locally semiperfect algebras $\Lambda$ over $R$ such that $\Lambda\otimes_{R}Q$ is simple artinian. In Section 6, we show that every trace of a projective module is the trace of a finitely generated projective module (cf. Lemma 6.6) and $V(\Lambda)\sqcup B(\Lambda)=V^{*}(\Lambda)$ (cf. Theorem 6.16). In Section 7, we study in detail the genus of projective modules over such algebras that we understand as a submonoid of the product of $V^{*}(\Lambda_{\mathfrak{m}})$ for $\mathfrak{m}$ maximal ideal of $R$ that can be described quite explicitly. In Proposition 7.1, we show that two countably generated projective modules that are not finitely generated are isomorphic if and only if all their localizations at maximal ideals of $R$ are isomorphic. That is, they are isomorphic if and only if they have the same genus.
@@
These tools allow us to study in Section 8 when every projective module is a direct sum of finitely generated ones. Finally, in Section 9, we apply our results to $\Lambda$ the endomorphism ring of a finitely generated torsion-free module $M$ over an $h$-local domain $R$ of Krull dimension $1$. In Theorem 9.3, we determine when any direct summand of an arbitrary direct sum of copies of $M$ is a sum of finitely generated modules. The results in this section complement the ones obtained in [AHP] in which we were studying domains for which the class of direct sums of finitely generated torsion-free modules is closed under direct summands.
@@ 1. The monoid of infinitely generated projective modules
For a given ring $\Lambda$, we write $V_{r}(\Lambda)$ (resp. $V_{r}^{*}(\Lambda)$) for the set of representatives of the isomorphism classes of finitely generated (resp. countably generated) projective right $\Lambda$-modules. If $P$ is a projective right $\Lambda$-module in $\mbox{\rm add}(\Lambda)$ (resp. in $\mbox{\rm Add}(\Lambda)$), we denote its representative in $V_{r}(\Lambda)$ (resp. in $V_{r}^{*}(\Lambda)$) by $[P]$.
@@
Unless stated otherwise, our modules will always be right modules, so we will write $V(\Lambda)$ and $V^{*}(\Lambda)$ instead of $V_{r}(\Lambda)$ and $V_{r}^{*}(\Lambda)$, respectively, and leave the previous notation in cases where both right and left modules appear. Every result stated for $V(\Lambda)$ and $V^{*}(\Lambda)$ can be stated in the context of left $\Lambda$-modules for $V_{\ell}(\Lambda)$ and $V_{\ell}^{*}(\Lambda)$, respectively.
@@
Note that both $V(\Lambda)$ and $V^{*}(\Lambda)$ are commutative monoids with addition defined by $[P]+[Q]=[P\oplus Q]$. Note that $V(\Lambda)$ is a submonoid of $V^{*}(\Lambda)$ and that $V^{*}(\Lambda)\setminus V(\Lambda)$ is a subsemigroup of $V^{*}(\Lambda)$.
An important object related to projective modules is trace ideals. We recall the definition and some basic properties.
@@
Alternatively, if $S=\mathrm{End}_{\Lambda}(M)$ the trace of $M$ is the image of the pairing
We will denote by $\mathcal{T}_{r}(\Lambda)$ (resp. $\mathcal{T}_{\ell}(\Lambda)$) the set of ideals of $\Lambda$ that are traces of countably generated projective right (resp. left) $\Lambda$-modules.
@@
${\tau\colon}$${\operatorname{Hom}_{\Lambda}(M,\Lambda)\otimes_{S}M}$${\Lambda}$${f\otimes x}$${f(x).}$
In the following, we will write $\mathcal{T}(\Lambda)$ instead of $\mathcal{T}_{r}(\Lambda)$, and leave the previous notation in cases where both right and left modules appear.
@@
Let $\Lambda$ be a ring, let $M,N$ be right $\Lambda$-modules, let $P$ be a projective right $\Lambda$-module, and let $J$ be an ideal of $\Lambda$. Then
@@
    If $M\subseteq_{\oplus}N$, then $\operatorname{Tr}_{\Lambda}(M)\subseteq\operatorname{Tr}_{\Lambda}(N)$.
@@
    $\operatorname{Tr}_{\Lambda}(M\oplus N)=\operatorname{Tr}_{\Lambda}(M)+\operatorname{Tr}_{\Lambda}(N)$.
@@
    [lam2, Proposition 2.40] $\operatorname{Tr}_{\Lambda}(P)\subseteq J$ if and only if $PJ=P$. In addition, the trace of a projective module is always an idempotent ideal.
@@
For a commutative ring $\Lambda$, Vasconcelos showed that trace ideals of projective modules are exactly the pure ideals.
In [whitehead], Whitehead gave a characterization of ideals that are traces of projectives for general rings. In particular, he showed that an idempotent ideal of $\Lambda$ that is finitely generated as a left ideal is always the trace of a countably generated projective right $\Lambda$-module. We shall use this result throughout the paper, sometimes without prior acknowledgment.
An alternative approach to these results characterizing the traces of projective modules can also be found in the paper [traces].
@@
Let $\Lambda$ be a ring with Jacobson radical $J(\Lambda)$. Let $I$ and $I^{\prime}$ be elements in $\mathcal{T}(\Lambda)$. Then $I=I^{\prime}$ if and only if $I+J(\Lambda)=I^{\prime}+J(\Lambda)$.
Therefore, if $I\in\mathcal{T}(\Lambda)$ then $I+J(\Lambda)/J(\Lambda)\in\mathcal{T}(\Lambda/J(\Lambda))$, and the assignment defined by $I\mapsto(I+J(\Lambda))/J(\Lambda)$ defines an injective map $\varphi\colon\mathcal{T}(\Lambda)\to\mathcal{T}(\Lambda/J(\Lambda))$.
Proof.  The first part of the statement is included in [traces, Corollary 2.9].
To prove the second part of the statement, note that if $I$ is the trace of the projective right $\Lambda$-module $P$ then, by Lemma 1.6, $I+J(\Lambda)/J(\Lambda)$ is the trace of $P/PJ(\Lambda)$. This shows that $\varphi$ is well defined. The injectivity of $\varphi$ follows from the first part of the statement.
The following lemma states that projective modules can be lifted modulo a trace ideal.

### v2

@@ Introduction
For a long time, the study of projective modules was focused on the finitely generated ones. Outside the finitely generated setting, the classical theorem of Kaplansky [kaplansky] stating that every projective module is a direct sum of countably generated projective modules shows that one can just consider the countably generated case. Interesting monoids associated to infinitely generated projective modules have recently been studied: the monoid $V(\Lambda)\sqcup B(\Lambda)$ of relatively big projectives defined, formally, in [wiegand] and that is the main object of interest in this paper, but also the analogous of the Cuntz monoid for general rings defined in [AABPV], and the $\kappa$-monoids defined in [nazemian_smertnig].
@@
The concept of relatively big projective module was introduced by Příhoda in [P2], extending Bass’ uniformly big projectives [bass]. Let $J$ be the trace of a countably generated projective module. A countably generated projective right $\Lambda$-module $P$ is said to be $J$-big, if every countably generated projective $\Lambda$-module whose trace is contained in $J$ is a homomorphic image of $P$. A countably generated projective module $P$ is said to be relatively $J$-big if $P$ is $J$-big and $P/PJ$ is finitely generated. Therefore, $J$ is maximal with respect to the property that $P$ is $J$-big and minimal with respect to the property that $P/PJ$ is finitely generated (cf. § 1.2). A crucial property is that a relatively $J$-big projective module $P$ is uniquely determined, up to isomorphism, by the trace ideal $J$ and the isomorphism class of $P/PJ$ (cf. Lemma 1.14).
The relatively big projectives are the countably generated projective modules that are relatively $J$-big for some $J$. The important fact that boils down in the concept of relatively big projective is that projective modules modulo a trace ideal can be lifted to projective modules over the ring. If the projective module over the quotient is finitely generated, then there is a unique lifting to a relatively big projective module over the ring.
@@
The monoid $V(\Lambda)\sqcup B(\Lambda)$ of isomorphism classes of relatively big projective modules was introduced in [wiegand] to capture the properties of such projectives. This monoid embeds in $V^{*}(\Lambda)$ and can be thought of as the disjoint union of $V(\Lambda/J)$, where $J$ runs through the trace ideals of the countably generated $\Lambda$-modules, see Corollary 2.5.
@@
In [P2], it was proved that over a noetherian ring $\Lambda$ which satisfies a certain descending condition on two-sided ideals, which we call the ()(*)-condition, $V(\Lambda)\sqcup B(\Lambda)=V^{*}(\Lambda)$ (cf. [P2, Lemma 2.8] and Theorem 1.23). Commutative noetherian rings, finitely generated algebras over commutative noetherian rings of Krull dimension one, and noetherian semilocal rings satisfy the ()(*)-condition. Therefore, all countably generated projective modules over these rings are relatively big. These results were the main ingredients used to characterize $V^{*}(\Lambda)$ for $\Lambda$ a noetherian semilocal ring in [crelle], or to study some infinite direct sum decompositions in [wiegand].
@@
The main results of this paper provide new classes of rings for which $V(\Lambda)\sqcup B(\Lambda)=V^{*}(\Lambda)$. We prove that right noetherian PI-rings satisfy the ()(*)-condition (see Theorem 4.8), and this is one of the main ingredients in showing that all countably generated right (or left) projective modules over a right noetherian PI-ring are relatively big in Corollary 4.9. In particular, this applies to module-finite algebras over commutative noetherian rings, such as group algebras of finite groups over commutative noetherian rings, or to endomorphism rings of finitely generated modules over commutative noetherian rings. This shows that the restriction on Krull dimension that we were imposing before was superfluous.
@@
Beyond the noetherian setting, it is easy to show that for semiperfect rings, all countably generated projective right (or left) modules are relatively big (cf. Proposition 6.4). In Theorem 6.17, we show that the same conclusion holds for $\Lambda$ a locally semiperfect torsion-free algebra, over an $h$-local commutative domain $R$ with field of fractions $Q$, such that $\Lambda\otimes_{R}Q$ is simple artinian.
@@
The $h$-local domains satisfy strong noetherianity conditions on the maximal spectrum. More precisely, any non-zero ideal of $R$ is contained only in a finite number of maximal ideals, and any non-zero prime ideal of $R$ is contained in a unique maximal ideal. Noetherian domains of Krull dimension one are always $h$-local. The definition of $h$-local domain is attributed to Matlis. However, Jaffard in [jaffard] introduced Dedekind-type domains as those domains $R$ such that any non-zero ideal is a finite product of ideals that are contained in a unique maximal ideal of $R$. Jaffard proved that Dedekind-type domains coincide with $h$-local domains, cf. [FHL, Theorem 2.1.5].
@@
The class of $h$-local domains is ubiquitously associated with many classical problems in the study of modules over non-noetherian domains, see [matlis3] and the recent survey [olberding]. Non-finitely generated projective modules over an $h$-local domain are free [hinohara]. Thus, what is of interest is the description of projective modules over algebras on such domains.
Our interest in this setting comes from our paper [AHP], in which we studied domains for which the class of direct sums of finitely generated torsion-free modules is closed under direct summands. If $\Lambda$ is the endomorphism ring of a finitely generated torsion-free module $M$ over an $h$-local domain $R$, then $\Lambda$ is torsion-free and $\Lambda\otimes_{R}Q$ is simple artinian. If, in addition, $R$ has Krull dimension $1$ and all the direct summands of a direct sum of copies of $R\oplus M$ are again a direct sum of finitely generated modules, it was proved in [AHP] that $\Lambda$ is locally semiperfect. Therefore, our results show that over such $\Lambda$ all countably generated projective modules are relatively big.
It is interesting to recall that Procesi and Small [PS] proved that the endomorphism ring of a finitely generated module over a commutative ring or, in general, over a ring satisfying a polynomial identity, is also a ring satisfying a polynomial identity. As far as we know, this point of view has not been so much examined in the world of finitely generated modules, endomorphism rings and direct sum decompositions, and certainly, it seems it should be worthwhile to explore it.
The paper is structured as follows. In Section 1 we give the basic definitions and notation. We introduce the needed background, presenting the concept of relatively big projective module, the ()(*)-condition and the main results relating both notions. In Section 2 we examine the monoid of isomorphism classes of relatively big projective modules, its algebraic properties and explore the right/left (non)symmetry of the monoids; check Example 2.7 and Corollary 2.9. In Section 3 we show that countably generated, projective right modules lift modulo a right $T$-nilpotent ideal, cf. Lemma 3.1. We prove that for a right $T$-nilpotent ideal $N$ of $\Lambda$ there are monoid isomorphisms between $V^{*}(\Lambda)$ and $V^{*}(\Lambda/N)$. This isomorphism restricts to isomorphisms between $V(\Lambda)$ and $V(\Lambda/N)$, and between $V(\Lambda)\sqcup B(\Lambda)$ and $V(\Lambda/N)\sqcup B(\Lambda/N)$, cf. Corollary 3.3. In Section 4, these results are an important ingredient to prove that right noetherian PI-rings satisfy the ()(*)-condition (Theorem 4.8), and to conclude that $V(\Lambda)\sqcup B(\Lambda)=V^{*}(\Lambda)$ for $\Lambda$ a right noetherian PI-ring in Corollary 4.9. In particular, this holds for module-finite algebras over commutative noetherian rings (cf. Corollary 4.10).
@@
In the rest of the paper, we turn to algebras over $h$-local domains. Section 5 discusses Package Deal Theorems for projective modules over $\Lambda$, whenever $\Lambda$ is an algebra over an $h$-local domain; our goal is twofold: recalling the results from [AHP] and proving some extensions needed in the subsequent sections. Section 6 focuses on semiperfect and locally semiperfect algebras. We observe in Proposition 6.4 that for semiperfect rings, all countably generated projective modules are relatively big.
@@
Let $R$ be an $h$-local domain with field of fractions $Q$. In the remaining sections, we study locally semiperfect algebras $\Lambda$ over $R$ such that $\Lambda\otimes_{R}Q$ is simple artinian. In Section 6, we show that for such algebras, every trace of a projective module is the trace of a finitely generated projective module (cf. Lemma 6.7) and that all countably generated projective $\Lambda$-modules are relatively big (cf. Theorem 6.17).
In Section 7, we study in detail the genus of countably generated projective modules over such algebras, that is, the isomorphism classes of countably generated projective modules over the localizations at the maximal ideals of $R$. This is a submonoid of the product of $V^{*}(\Lambda_{\mathfrak{m}})$, for $\mathfrak{m}$ maximal ideal of $R$, that can be described quite explicitly in terms of solutions in ${\mathbb{N}}_{0}\cup\{\infty\}$ of a system of linear equations with coefficients in ${\mathbb{N}}_{0}$.
In Proposition 7.1, we show that two countably generated projective modules that are not finitely generated are isomorphic if and only if they have the same genus, that is, whenever all their localizations at maximal ideals of $R$ are isomorphic. This is also true for finitely generated projective modules, provided $\Lambda$ is a finitely generated algebra over a semilocal domain $R$, cf. Proposition 7.2.
@@
These tools allow us to study in Section 8 when every projective module is a direct sum of finitely generated modules. Finally, in Section 9, we apply our results to $\Lambda$ the endomorphism ring of a finitely generated torsion-free module $M$ over an $h$-local domain $R$ of Krull dimension $1$. In Theorem 9.3, we give necessary and sufficient conditions that imply that any direct summand of an arbitrary direct sum of copies of $M$ is a sum of finitely generated modules. These results complement and clarify the ones obtained in [AHP].
@@
The set $\mathbb{N}=\{1,2,\dots\}$ and ${\mathbb{N}}_{0}=\mathbb{N}\cup\{0\}$.
We thank the reviewer for the careful reading of the article, and for all the helpful comments and suggestions.
@@ 1. The monoid of infinitely generated projective modules
For a given ring $\Lambda$, we write $V_{r}(\Lambda)$ (resp. $V_{r}^{*}(\Lambda)$) for the set of representatives of the isomorphism classes of finitely generated (resp. countably generated) projective right $\Lambda$-modules. If $P$ is a finitely (resp. countably) generated projective right $\Lambda$-module, we denote its representative in $V_{r}(\Lambda)$ (resp. in $V_{r}^{*}(\Lambda)$) by $[P]$.
@@
Unless stated otherwise, our modules will always be right modules, so we will write $V(\Lambda)$ and $V^{*}(\Lambda)$ instead of $V_{r}(\Lambda)$ and $V_{r}^{*}(\Lambda)$, respectively, and leave the previous notation in cases where both right and left modules appear. All the results stated for $V(\Lambda)$ and $V^{*}(\Lambda)$ can be stated in the context of left $\Lambda$-modules for $V_{\ell}(\Lambda)$ and $V_{\ell}^{*}(\Lambda)$, respectively.
@@
Note that both $V(\Lambda)$ and $V^{*}(\Lambda)$ are commutative monoids with addition defined by $P]+[Q]=[P\oplus Q]$. Note that $V(\Lambda)$ is a submonoid of $V^{*}(\Lambda)$ and that $V^{*}(\Lambda)\setminus V(\Lambda)$ is a subsemigroup of $V^{*}(\Lambda)$.
An important object related to projective modules is the set of their trace ideals. We recall the definition and some basic properties.
@@
We will denote by $\mathcal{T}_{r}(\Lambda)$ (resp. $\mathcal{T}_{\ell}(\Lambda)$) the set of ideals of $\Lambda$ that are traces of countably generated projective right (resp. left) $\Lambda$-modules. When we simply write $\mathcal{T}(\Lambda)$ we mean $\mathcal{T}_{r}(\Lambda)$.
@@
$\tau\colon$$\operatorname{Hom}_{\Lambda}(M,\Lambda)\otimes_{S}M$$\Lambda$$f\otimes x$}f(x).{f(x).}
@@
Let $\Lambda$ be a ring.
@@
    Let $M,N$ be right $\Lambda$-modules. If $M$ is a homomorphic image of $N$, then $\operatorname{Tr}_{\Lambda}(M)\subseteq\operatorname{Tr}_{\Lambda}(N)$.
@@
    Let $\{M_{i}\}_{i\in A}$ be a family of right $\Lambda$-modules. Then
    $\operatorname{Tr}_{\Lambda}\Big(\bigoplus_{i\in A}M_{i}\Big)=\sum_{i\in A}\operatorname{Tr}_{\Lambda}(M_{i}).$
@@
    [lam2, Proposition 2.40] Let $J$ be an ideal of $\Lambda$, and let $P$ be a projective right $\Lambda$-module. Then $\operatorname{Tr}_{\Lambda}(P)\subseteq J$ if and only if $PJ=P$. In addition, the trace of a projective module is always an idempotent ideal.
@@
For a commutative ring $\Lambda$, Vasconcelos showed that trace ideals of projective modules are exactly the pure ideals (see page 430 in [Vasconcelos2] or Proposition 2.12 in [traces]).
In [whitehead], Whitehead gave a characterization of ideals that are traces of projective modules for general rings. In particular, he showed that an idempotent ideal of $\Lambda$ that is finitely generated as a left ideal is always the trace of a countably generated projective right $\Lambda$-module. We shall use this result throughout the paper, sometimes without prior acknowledgment.
An alternative approach to the characterization of the traces of projective modules can be found in the paper [traces].
@@
Let $\Lambda$ be a ring with Jacobson radical $J(\Lambda)$, and let $I$ and $I^{\prime}$ be elements in $\mathcal{T}(\Lambda)$. Then
  1. (i)
    $I=I^{\prime}$ if and only if $I+J(\Lambda)=I^{\prime}+J(\Lambda)$.
  2. (ii)
    $(I+J(\Lambda))/J(\Lambda)\in\mathcal{T}(\Lambda/J(\Lambda))$, so the assignment defined by $I\mapsto(I+J(\Lambda))/J(\Lambda)$ defines an injective map $\varphi\colon\mathcal{T}(\Lambda)\longrightarrow\mathcal{T}(\Lambda/J(\Lambda))$.
Proof. $(i)$ is included in [traces, Corollary 2.9].
To prove $(ii)$, note that if $I$ is the trace of the projective right $\Lambda$-module $P$ then, by Lemma 1.6, $(I+J(\Lambda))/J(\Lambda)$ is the trace of $P/PJ(\Lambda)$. This shows that $\varphi$ is well defined. The injectivity of $\varphi$ follows from $(i)$.
The following lemma states that projective modules can be lifted modulo a trace ideal. This is a crucial result in the development of the theory of relatively big projective modules.